# NVRC: Neural Video Representation Compression

### v1

@@ 5 Conclusion
In this paper, we present NVRC, a new INR-based video compression framework with a focus on representation compression. By employing novel entropy coding and quantization models, NVRC significantly improved coding efficiency and allows real end-to-end optimization for the INR model. The experimental results show that NVRC outperforms all the benchmarked conventional and learning-based video codecs, in particular with a 24% bitrate saving against VVC VTM (Random Access) on the UVG database. This is the first time an INR-based video codec has obtained this achievement. Future works should focus on reducing the complexity of this model.
@@
Appendix A Appendix / Supplemental material
A.1 NVRC configurations at different scales.
Neural representation. In NVRC, the INR architecture is based on HiNeRV [25], with minor modifications. The differences include: (1) due to the enhanced capability of feature grids coding, feature grids in NVRC are larger for better capability on capturing dynamic contents. (2) The bilinear interpolation of the input encoding is performed after the stem convolution layer, while in the original HiNeRV it was before applying convolution. This improves the performance marginally. (3.) The hyper-parameter selection between different scales has been simplified, where only the number of channels of the network layers and the feature grids change with scales.
The hyper-parameters of the neural representation in NVRC are provided in Table 5 and 6. It is noted that in NVRC, only the input feature grids in HiNeRV are coded by the context model, while the local grids are simply considered as general layer parameters, as they only account for a small number of parameters.
@@ Appendix A Appendix / Supplemental material
Zhang et al. [59] D R+D N/A Uniform per-channel No
Gomes et al. [14] D R+D N/A Neural Network per-weight No
Maiya et al. [38] R+D R+D N/A Neural Network per-weight No
Kim et al. [24] R+D R+D Context Laplace* No No
Leguay et al. [28] R+D - Context Laplace* No No
Zhang et al. [58] D R+D Gaussian Gaussian per-weight No
@@
Grid sizes $200\times 45\times 80\times 1$ $200\times 45\times 80\times 2$ $200\times 45\times 80\times 4$ $200\times 45\times 80\times 8$
@@
Context-based entropy model. The context-based entropy model in NVRC has an autoregressive style coding process. It is based on masked convolution [52, 42]. It contains 3 blocks, where each block contains a Layer Normalization layer [4], 3D convolution and GeLU activation [19] (except for the final output). The context-based model is independent between channels, and has been implemented as with depth-wise convolution. The kernel size is 5 and the width is 8. The output of the context-based model is the means and log scales of the Gaussian distribution.
@@
A.2 Experiment configurations
In the experiments, the configurations of NVRC are based on [25]. The training is performed by sampling patches from the target video, where the patch size is $120\times 120$, and the batch size of each training step is 144 patches (equal to 1 frame). For learning RGB output, the distortion loss is:
@@ A.2 Experiment configurations
The learning rates in Stage 1 and Stage 2 are 2e-3 and 1e-4, where the cosine decay is applied for scaling the learning rate with a minimum learning rate of 1e-4 and 1e-5, respectively. The norm clipping with 1.0 is applied. L2 regularization of 1e-6 is applied to improve the numerical stability, as we observe that the norm of weight could grow too large in some cases. The magnitude of L2 regularization linearly decays in the first stage and is not applied in the second stage, to avoid under-fitting. For the soft-rounding and the Kumaraswamy noise [24] in Stage 1, the temperatures and the noise scale ratio scale from 0.5 to 0.3 and 2.0 to 1.0, respectively. For Quant-Noise [48] in Stage 2, the noise ratio scales from 0.5 to 1.0. Note that we follow [25], where rounding, instead of the straight-through estimator (STE) [7], is applied to the quantized variables. $R$ is optimized once for every 8 steps of $D$.
Table 6: NVRC representation scales and $\lambda$ configurations for the UVG datasets (RGB/YUV).
Rate point UVG (RGB) UVG (YUV)
1 S2, $\lambda=1$ S2, $\lambda=32$
2 S2, $\lambda=2$ S2, $\lambda=64$
3 S3, $\lambda=4$ S3, $\lambda=128$
4 S3, $\lambda=8$ S3, $\lambda=256$
5 S4, $\lambda=16$ S4, $\lambda=512$
6 S4, $\lambda=32$ S4, $\lambda=1024$
The results for HiNeRV [25] and HNeRV-Boost [58] are obtained by training with their provided implementations, with the original configurations. The results of C3 [24] in the UVG dataset is obtained from the original paper.
@@ 6 S4, $\lambda=32$ S4, $\lambda=1024$
A.2.1 Rate distribution
In Fig 4 (left), we provide the rate distribution of different types of parameters (feature grids, network layer parameters, quantization and entropy coding parameters for feature grids/layer parameters). The data is collected from the lowest rate point models with the UVG dataset [40]. It can be observed that, in general, the quantization and entropy coding parameters contribute to a very small amount of the total bitrate, where the ratio between the feature grids and the layer parameters vary between different video sequences. In Figure 4 (right), we further provide the bits-per-parameter data for different video sequences. In this very low bitrate rate point, NVRC is capable of learning parameter distributions with very low bits-per-parameters, and it also varies between sequences. For example, on some sequences with larger motion (Jockey and YachtRid), the bits-per-parameters of the feature grids can be doubled, but for some relatively static sequences, the bits-per-parameters of the feature grids is nearly zero.
A.3 Positive Impacts
NVRC is the first INR-based codec which outperforms VVC VTM with a 24% coding gain on the UVG database. It will potentially contribute to the next generation of video coding standards, and improve current video streaming services if deployed in practice.
A.4 Limitations
As the implementation of NVRC is based on sophisticated neural networks, it requires substantial computational resources, in particular at the encoder. This results in increased energy consumption and a negative environmental impact. Future work should focus on reducing the complexity of this model.

### v2

@@ 5 Conclusion
In this paper, we present NVRC, a new INR-based video compression framework with a focus on representation compression. By employing novel entropy coding and quantization models, NVRC significantly improved coding efficiency and allows real end-to-end optimization for the INR model. The experimental results show that NVRC outperforms all the benchmarked conventional and learning-based video codecs, in particular with a 23% bitrate saving against VVC VTM (Random Access) vtm on the UVG database mercat2020uvg. This is the first time an INR-based video codec has obtained this achievement.
@@
Appendix A Appendix
A.1 Additional experiments
Table 5: BD-rate results on the JVET-CTC Class B dataset boyce2018jvet.
Color Space Metric x265 (veryslow) HM (RA) VTM (RA) DCVC-DC
RGB 4:4:4 PSNR -68.70% -35.53% 7.48% -13.85%
MS-SSIM -84.69% -65.51% -42.60% -18.81%
YUV 4:2:0 PSNR -66.33% -34.49% 9.57% -24.25%
MS-SSIM -66.02% -39.99% -6.79% -36.94%
Figure 5: Average rate quality curves of various tested codecs on the JVET-CTC Class B datasets boyce2018jvet.
In addition to the main paper experiments, here we provide the results of NVRC with the JVET-CTC test sequences for VTM boyce2018jvet. Similar experiments configurations as in the main paper have been used, where x265 x265 with the configuration, HM-18.0 hm and VTM-20.0 vtm with the Random Access configuration, and DCVC-DC li2023neural have been used as the baseline models. Similar to the primary experiment results, our proposed NVRC consistently outperforms the baseline models in most cases.
A.2 NVRC configurations at different scales.
Neural representation. In NVRC, the INR architecture is based on HiNeRV kwan2023hinerv, with minor modifications. The differences include: (1) due to the enhanced capability of feature grids coding, feature grids in NVRC are larger for better capability on capturing dynamic contents. (2) The bilinear interpolation of the input encoding is performed after the stem convolution layer, while in the original HiNeRV it was before applying convolution. This improves the performance marginally. (3.) The hyper-parameter selection between different scales has been simplified, where only the number of channels of the network layers and the feature grids change with scales.
The hyper-parameters of the neural representation in NVRC are provided in Table 7 and 8. It is noted that in NVRC, only the input feature grids in HiNeRV are coded by the context model, while the local grids are simply considered as general layer parameters, as they only account for a small number of parameters.
@@ A.1 Additional experiments
Zhang et al. zhang2021implicit D R+D N/A Uniform per-channel No
Gomes et al. gomes2023video D R+D N/A Neural Network per-weight No
Maiya et al. maiya2023nirvana R+D R+D N/A Neural Network per-weight No
Kim et al. kim2023c3 R+D R+D Context Laplace* No No
Leguay et al. leguay2024cool R+D - Context Laplace* No No
Zhang et al. zhang2024boosting D R+D Gaussian Gaussian per-weight No
@@
Grid sizes $T_{grid}\times 45\times 80\times 1$ $T_{grid}\times 45\times 80\times 2$ $T_{grid}\times 45\times 80\times 4$ $T_{grid}\times 45\times 80\times 8$
@@
$T_{grid}$: 200 for UVG mercat2020uvg/JVET-CTC Class B boyce2018jvet, 50 for MCL-JCV wang2016mcl
Context-based entropy model. The context-based entropy model in NVRC has an autoregressive style coding process. It is based on masked convolution oord16conditional; minnen18joint. It contains 3 blocks, where each block contains a Layer Normalization layer ba2016layer, 3D convolution and GeLU activation hendrycks16bridging (except for the final output). The context-based model is independent between channels, and has been implemented as with depth-wise convolution. The kernel size is 5 and the width is 8. The output of the context-based model is the means and log scales of the Gaussian distribution.
@@
A.3 Experiment configurations
In the experiments, the configurations of NVRC are based on kwan2023hinerv. The training is performed by sampling patches from the target video, where the patch size is $120\times 120$, and the batch size of each training step is 144 patches (equal to 1 frame). For learning RGB output, the distortion loss is:
@@ A.3 Experiment configurations
The learning rates in Stage 1 and Stage 2 are 2e-3 (or 1e-3 for rare case which the training is less stable) and 1e-4, where the cosine decay is applied for scaling the learning rate with a minimum learning rate of 1e-4 and 1e-5, respectively. The norm clipping with 1.0 is applied. L2 regularization of 1e-6 is applied to improve the numerical stability, as we observe that the norm of weight could grow too large in some cases. The magnitude of L2 regularization linearly decays in the first stage and is not applied in the second stage, to avoid under-fitting. For the soft-rounding and the Kumaraswamy noise kim2023c3 in Stage 1, the temperatures and the noise scale ratio scale from 0.5 to 0.3 and 2.0 to 1.0, respectively. For Quant-Noise fan2020training in Stage 2, the noise ratio scales from 0.5 to 1.0. Note that we follow kwan2023hinerv, where rounding, instead of the straight-through estimator (STE) bengio13estimating, is applied to the quantized variables. $R$ is optimized once for every 8 steps of $D$.
Table 8: NVRC representation scales and regularization ($\lambda$) configurations for the UVG mercat2020uvg, MCL-JCV wang2016mcl and JVET-CTC Class B boyce2018jvet datasets (RGB/YUV).
Rate point UVG MCL-JCV JVET-CTC Class B
RGB YUV RGB YUV RGB YUV
1 S2, $\lambda=1.0$ S2, $\lambda=32.0$ S1, $\lambda=0.25$ S1, $\lambda=8.0$ S2, $\lambda=1.0$ S2, $\lambda=32.0$
2 S2, $\lambda=2.0$ S2, $\lambda=64.0$ S1, $\lambda=0.5$ S1, $\lambda=16.0$ S2, $\lambda=2.0$ S2, $\lambda=64.0$
3 S3, $\lambda=4.0$ S3, $\lambda=128.0$ S2, $\lambda=1.0$ S2, $\lambda=32.0$ S3, $\lambda=4.0$ S3, $\lambda=128.0$
4 S3, $\lambda=8.0$ S3, $\lambda=256.0$ S2, $\lambda=2.0$ S2, $\lambda=64.0$ S3, $\lambda=8.0$ S3, $\lambda=256.0$
5 S4, $\lambda=16.0$ S4, $\lambda=512.0$ S3, $\lambda=4.0$ S3, $\lambda=128.0$ S4, $\lambda=16.0$ S4, $\lambda=512.0$
6 S4, $\lambda=32.0$ S4, $\lambda=1024.0$ S3, $\lambda=8.0$ S3, $\lambda=256.0$ S4, $\lambda=32.0$ S4, $\lambda=1024.0$
For the conventional codecs, we used a QP range of 16–44 for x265 x265 and 16–36 for HM hm and VTM vtm, with QP values selected at intervals of 4. The results for HiNeRV kwan2023hinerv and HNeRV-Boost zhang2024boosting are obtained by training with their provided implementations, with the original configurations.
@@
A.3.1 Rate distribution
In Fig 6 (left), we provide the rate distribution of different types of parameters (feature grids, network layer parameters, quantization and entropy coding parameters for feature grids/layer parameters). The data is collected from the lowest rate point models with the UVG dataset mercat2020uvg. It can be observed that, in general, the quantization and entropy coding parameters contribute to a very small amount of the total bitrate, where the ratio between the feature grids and the layer parameters vary between different video sequences. In Figure 6 (right), we further provide the bits-per-parameter data for different video sequences. In this very low bitrate rate point, NVRC is capable of learning parameter distributions with very low bits-per-parameters, and it also varies between sequences. For example, on some sequences with larger motion (Jockey and YachtRide), the bits-per-parameters of the feature grids can be doubled, but for some relatively static sequences, the bits-per-parameters of the feature grids is nearly zero.
A.4 Positive Impacts
NVRC is the first INR-based codec which outperforms VVC VTM vtm with a 23% coding gain on the UVG database mercat2020uvg. It will potentially contribute to the next generation of video coding standards, and improve current video streaming services if deployed in practice.
A.5 Limitations
Encoder complexity. As the implementation of NVRC is based on sophisticated neural networks, it requires substantial computational resources, in particular at the encoder. This is a common issue with many INR-based approaches that require content overfitting during encoding. It makes this type of approaches unsuitable for real-time encoding scenarios like video conferencing. This also results in increased energy consumption and a negative environmental impact. Future work should focus on reducing the complexity of this model.
Latency. As INR-based video codecs require overseeing all frames of a video sequence at the same time during encoding, the system latency become more longer compared to conventional and some end-to-end learned video codecs which perform per-frame encoding. This prevents these INR models from adoption in practical application scenarios.
Reproducibility. In this paper, we have not studied reproducibility, which is a critical issue for the practical application of deep video compression with entropy coding. While our experiments focus on floating-point operations, the operations in the proposed method can also be implemented, for example, using integer operations ballejm19, which can ensure reproducibility.